OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin^2 x - cos^2 x = 0

OpenStudy (anonymous):

Tried setting sin^2 x and -cos^2 x equal to zero, but that didn't work out... Answer choices: x = pi divided by four x = pi divided by four, three pi divided by four, five pi divided by four, seven pi divided by four x = seven pi divided by four, pi divided by three x = seven pi divided by four, pi divided by six

OpenStudy (tamtoan):

you know that : \[\sin^{2} x + \cos^{2} x = 1 \] right? apply this to see what you get.

OpenStudy (anonymous):

I don't really understand what you mean.

OpenStudy (tamtoan):

with the formula there, you can eliminate either sin^2 (x) or cos^2(x) ...then solve it.

OpenStudy (anonymous):

sin^2 x = 0?

OpenStudy (anonymous):

not zero, 1?

OpenStudy (tamtoan):

let's add the 2 equations you have there and see what you get? ...sin^2(x) is not 1 ..it's twice of that is 1

OpenStudy (anonymous):

2 sin^2 x = 1?

OpenStudy (tamtoan):

yeah , then solve for x :)

OpenStudy (anonymous):

x = pi divided by four, three pi divided by four, five pi divided by four, seven pi divided by four ?

OpenStudy (sirm3d):

\[\Huge \color{blue} \checkmark \]

OpenStudy (tamtoan):

well, you have sin^2(x) = 1/2 ...take square root both side and you have sin(x) = \[1/\sqrt{2}\] from here, you can use calculator to find out...i don't remember which one.